\begin{thebibliography}{ppp}


\bibitem[AD]{ad} S.A. Argyros,  I. Deliyanni, \textit{Examples of asymptotic $\ell_1$ Banach Spaces}, Trans. Amer. Math. Soc. 349 (1997), 973--995.

\bibitem[AT]{at} S.A. Argyros,  S. Todorcevic, \textit{Ramsey Methods in Analysis}, Advanced Course in Mathematics CRM Barcelona,
Birkh\"auser Verlag, Basel-Boston-Berlin, 2005.

\bibitem[BL]{bl} S.J. Bernau, H.E. Lacey, \textit{The range of a contractive projection on an $L_p$-space},
Pacific J. Math. 53 (1974), 21-41.

\bibitem[CFS]{cfs} S. Campbell, G. Faulkner,  R. Sine, \textit{Isometries, projections and Wold decompositions},  in Operator theory and functional analysis (Papers, Summer Meeting, Amer. Math. Soc., Providence, R.I., 1978), Pitman, Boston, Mass., 1979, 85–114. 

\bibitem[D]{d} S. Ditor, \textit{Averaging operators in $C(S)$ and lower semicontinuous sections of continuous maps}, Trans. Amer. Math. Soc., 175 (1973), 195–208.


\bibitem[FH]{fh} G.D. Faulkner, J.E. Huneycutt, Jr., \textit{Orthogonal decomposition of isometries in a Banach space}, Proc. Amer. Math. Soc., 69 (1978), 125–128.


\bibitem[FJ]{fj} R. J. Fleming, J. E. Jamison, \textit{Isometries on Banach spaces: function spaces}, Chapman \& Hall/CRC Monographs and Surveys in Pure and Applied Mathematics, 129,
Chapman \& Hall/CRC, Boca Raton, FL, 2003.

\bibitem[GM1]{gm} W.T. Gowers, B. Maurey, \textit{The unconditional basic sequence problem}, J. Amer. Math. Soc., 6 (1993), no. 4, 851--874.

\bibitem[GM2]{gm2} W.T. Gowers,  B. Maurey, \textit{Banach spaces with small spaces of operators},  Math. Ann.  307  (1997),  no. 4, 543--568.

\bibitem[LT]{lt} J. Lindenstrauss, L. Tzafriri, \textit{Classical Banach Spaces I. Sequence Spaces}, Springer- Verlag, Berlin-Heidelberg-New York 1977. %Prop. 1.a.9

\bibitem[M]{m} B. Maurey, \textit{Banach spaces with few operators}.  Handbook of the geometry of Banach spaces, Vol. 2,  1247--1297, NortHolland, Amsterdam, 2003.

\bibitem[Mb]{mb} M. Mbekhta, \textit{Partial isometries and generalized inverses}, Acta Sci. Math. (Szeged) 70 (2004), 767-781.

\bibitem[P]{p} A. Pe{\l}czy\'nski, \textit{Projections in certain Banach spaces},  Studia Math. 19 (1960), 209-228.

\bibitem[R]{r} B. Randrianantoanina, \textit{Norm-one projections in Banach spaces}, Taiwan. J. Math. 5 (2001), 35-95.

\bibitem[S]{s} T. Schlumprecht, \textit{An arbitrary distortable Banach space}, Israel J. Math. 76 (1991),  81--95.

\bibitem[T]{t} L. Tzafriri, \textit{Remarks on contractive projections in $L_p$-spaces}, Israel J. Math. 7 (1969), 9-15.

\end{thebibliography}

\documentclass[11pt]{amsart}
\usepackage[margin=2.5cm]{geometry}
\usepackage{amssymb}
\usepackage{bm}
\usepackage[centertags]{amsmath}
\usepackage{amsfonts}
\usepackage{amsthm}
\usepackage{mathrsfs}
\usepackage{mathtools,xfrac}
\usepackage[usenames,dvipsnames,svgnames,table]{xcolor}
\usepackage[mathscr]{euscript}
\usepackage[shortlabels]{enumitem}

\DeclareMathOperator\w{w}
\DeclareMathOperator\range{rng}

\DeclareMathOperator\supp{supp}

\DeclareMathOperator\suc{succ}

\newcommand{\mt}{\mathcal{T}}

\numberwithin{equation}{section}

\newtheorem{theorem}{Theorem}[section]
\newtheorem*{theorem*}{Theorem}
\newtheorem{lemma}[theorem]{Lemma}
\newtheorem{proposition}[theorem]{Proposition}
\newtheorem{corollary}[theorem]{Corollary}

\newtheorem{fact}[theorem]{Fact}

\newtheorem{definition}[theorem]{Definition}

\newtheorem{assumption}[theorem]{Assumption}

\newtheorem{remark}[theorem]{Remark}

\newcommand{\N}{\mathbb{N}}
\newcommand{\R}{\mathbb{R}}

\usepackage[utf8]{inputenc}

\author{Anna Pelczar-Barwacz}
\thanks{The research of the author was supported by the grant of the National Science Centre (NCN), Poland, no. UMO-2020/39/B/ST1/01042}

\email{anna.pelczar@uj.edu.pl}
\address{Jagiellonian University, Faculty of Mathematics and Computer Science, Institute of Mathematics,  {\L}ojasiewicza 6, 30-348 Krak\'ow,  Poland}

\usepackage[T1]{fontenc}
\numberwithin{subsection}{section}
\numberwithin{equation}{section}

\begin{document}
\title{A reflexive Banach space with an uncomplemented isometric subspace} 

\begin{abstract}
We construct a reflexive Banach space $X$ with a subspace isometric to $X$, which is not complemented in $X$. 
\end{abstract}
\maketitle

\section*{Introduction}

The study of the family of isometries on Banach spaces usually concentrates on the group of surjective isometries. Among non-surjective isometries, the class with particularly regular properties, parallel to those of isometries on Hilbert spaces, consists of isometries with 1-complemented range, that is with range equal to the image of some norm-one projection. The most useful result concerns the Wold Decomposition, well known in the case of Hilbert spaces, facilitating further description of isometries. Any isometry on a reflexive Banach space with 1-complemented range is a "Wold isometry", i.e. satisfies an extended version of the Wold decomposition, proved in the case of smooth reflexive spaces in \cite{fh} and for all reflexive spaces in \cite{cfs}. Note that the reverse implication does not hold true in general, as there is a Wold isometry on $C(K)$-space with complemented, but not 1-complemented range \cite{cfs}. we recall also that, in the case of complex Banach spaces, an isometry has 1-complemented range iff it admits a contractive generalized inverse, a property that characterizes partial isometries on Hilbert spaces by \cite{mb}.  

These results give rise to the question on isometries on Banach spaces with range that is not 1-complemented, first formulated in \cite{fj}, see also \cite{cfs} and \cite[Section 5.g]{r}. For results recalled below and related see an excellent survey \cite{r}, for the general study of isometries - \cite{fj}. 

It is well-known that all isometries in Hilbert spaces have 1-complemented ranges. The same property holds in $L_p$-spaces, $1\leq p<\infty$, on any measure space \cite{t}, see also \cite{bl},  in the case of $\ell_p$ spaces proved first in \cite{p}. On the other hand, $C[0,1]$ contains an isometric subspace that is not complemented \cite{d}, which leaves open the reflexive case.  
We present the following result which answers  the above question.

\begin{theorem}\label{main}
There is a reflexive Banach space $X$ with a subspace isometric to $X$ and not complemented in $X$.

More precisely, $X$ has a basis $(e_i)_{i=1}^\infty$ such that the mapping carrying each basic vector $e_i$ to $e_{2i}$, $i\in\N$, extends to an isometry between $X$ and $Y:=\overline{\textrm{span}}\{e_{2i}: i\in\N\}$ with $Y$ not complemented in $X$.   
\end{theorem}

The natural setting for search for a space $X$ satisfying the above theorem 
is the class of Banach spaces with strictly limited family of complemented subspaces, more general: with small algebra of bounded operators, namely gothe spaces of Gowers-Maurey type. The Gowers-Maurey technique, started in \cite{gm} and developed in \cite{gm2}, see also \cite{m}, yields Banach spaces with prescribed  algebra of bounded operators, including the first known HI (hereditarily indecomposable) Banach space \cite{gm}, constructed as a solution to the unconditional basis sequence problem. Recall that a HI space is a space with no decomposable subspaces, where a space is decomposable if it can be written as a direct sum of two its infinite dimensional subspaces. The theory of spaces with small algebra of operators, with its origins in the construction Tsirelson space, the first known space containing no copies of $\ell_p$, $1\leq p<\infty$, or $c_0$, developed extensively in last years, bringing the celebrated Argyros-Haydon space with "scalar+compact" property. 

We recall the general result of \cite{gm2}, as it forms a starting point for the construction presented here. Given two infinite $A=(a_i)_{i=1}^\infty,B=(b_i)_{i=1}^\infty\subset\N$,  a spread $S_{A,B}$ is a linear mapping $S_{A,B}:c_{00}\to c_{00}$ carrying each basic vector $e_{a_i}$ to $e_{b_i}$, $i\in\N$, and each basic vector $e_i$, $i\not\in A$, to 0. A family of spreads $\mathcal{S}$  is called proper, if it is closed under composition and taking adjoints (i.e. for any $S_{A,B}\in\mathcal{S}$ also $S_{B,A}\in\mathcal{S}$), and not "too big"  in a certain sense, in particular is countable. Note that if $\mathcal{S}$ is proper, for any $S_{A,B}\in\mathcal{S}$ also $S_{A,A}$, a projection onto $\textrm{span}\{e_i: i\in A\}$, is in $\mathcal{S}$. In \cite{gm2} for each proper family of spreads $\mathcal{S}$ a Banach space $X(\mathcal{S})$ is constructed as a completion of $c_{00}$ under certain norm, such that each element  $S_{A,B}\in\mathcal{S}$ extends to a partial isometry (carrying isometrically $\textrm{span}\{e_i:i\in A\}$ to $\textrm{span}\{e_i:i\in B\}$, with $\overline{\textrm{span}}\{e_i:i\in A\}$ and $\overline{\textrm{span}}\{e_i:i\in B\}$ 1-complemented in $X(\mathcal{S})$). What is crucial, the Gowers-Maurey tools guarantee that, up to strictly singular perturbation, the algebra of bounded operators on $X(\mathcal{S})$ is generated by $\mathcal{S}$. 

Taking for instance $\mathcal{S}$  to be a proper family of spreads generated by $S_{\N,2\N}$ we obtain a space $X(\mathcal{S})$ isometric to its subspace $Y=\overline{\textrm{span}}\{e_{2i}: i\in\N\}$, which is 1-complemented in $X$. We shall perturb the Gowers-Maurey construction, keeping the spread $S_{\N,2\N}$ as an isometry, but eliminating the continuity of the  adjoint spread $S_{2\N,\N}$ and projection $S_{2\N,2\N}$, by extensive use of Gowers-Maurey tools. Concerning the methodology, instead of following directly  \cite{gm,gm2} we built the space $X$ within the classical by now framework of mixed Tsirelson spaces that was initiated in \cite{ad} and discussed in great detail in \cite{at}.

The paper is organized as follows. In Section 1 we set the standard terminology, including mixed Tsirelson spaces and present the construction of the space $X$, which yields immediately an isometry between $X$ and its subspace $Y$ (Prop. \ref{isometric}). We recall the classical Gowers-Maurey tools following directly \cite{at} in Section 2,  concluding it with the result on reflexivity of $X$ (Corollary \ref{reflexive}). Section 3 is devoted to the proof of the fact, that $Y$ is not complemented in $X$ (Theorem \ref{not-complemented}) based again on an extensive use of the Gowers-Maurey technique. 

\section{The definition of the space $X$}

We set first the standard notation. We let
$\N$ denote the set of positive integers and $\N_0=\N\cup\{0\}$. Given any interval $E=[r,s]\subset\N$ and $k\in\N$ write $kE=[kr,ks]$.  $c_{00}$ denote the vector space of sequences in $\R$ which are eventually  zero. We shall treat the elements of $c_{00}$ either as vectors, denoted by $x,y,\dots$, with the unit vector basis  $(e_i)_{i=1}^\infty$, or as functionals, denoted by $f,g,\dots$, with the biorthogonal basis $(e_i^*)_{i=1}^\infty$.  
The {support} of a vector $v=(a_i)_i\in c_{00}$ is the set $\supp (v)=\{i\in\N: a_i\neq 0\}$, $\range(v)$ - the smallest interval in $\N$ containing the support of $x$. For any
$v=(a_i)_i\in c_{00}$ and $E\subset\N$ let $Ev\in c_{00}$ be defined by $(Ev)_i=a_i$ if $i\in E$ and $(Ev)_i=0$ otherwise.  
We write $v<u$ for  $v,w\in c_{00}$, if $\max\supp (v)<\min \supp (w)$. A block sequence is any  sequence $(v_n)_n\subset c_{00}$ satisfying $v_{1}<v_{2}<\dots$, a block subspace of $c_{00}$ - any  subspace spanned by an infinite block sequence. 

\subsection{Mixed Tsirelson spaces} Spaces of Gowers-Maurey type are built on the basis of  mixed Tsirelson spaces, the theory of which started with the Schlumprecht space \cite{s}, the first known arbitrary distortable space. We recall now the definition and basic properties of mixed Tsirelson spaces, defined by low complexity families $(\mathcal{A}_n)_{n=1}^\infty$, and spaces built on their basis, following \cite{at}. 

Fix two increasing sequences $(n_j)_{j=1}^\infty,(m_j)_{j=1}^\infty\subset\N$ with $m_j<n_j$, $j\in\N$. 

\begin{definition}[Mixed Tsirelson space]  Let $\tilde{K}\subset c_{00}$ be the smallest set satisfying the following:
\begin{enumerate}
    \item $(\pm e_i^*)_{i=1}^\infty\subset \tilde{K}$,
    \item for any $j\in\N$ and $f_1<\dots<f_d$ in $\tilde{K}$ with $d\leq n_j$ we have $m_j^{-1}(f_1+\dots+f_d)\in \tilde{K}$.
\end{enumerate}
Let $\|\cdot\|$ be the norm on $c_{00}$ with the norming set $\tilde{K}$, i.e. $\|\cdot\|=\sup\{|f(\cdot)|:f\in \tilde{K}\}$. The mixed Tsirelson space
$T[(\mathcal{A}_{n_j},m_j^{-1})_{j\in\N}]$ is the completion of $(c_{00}, \|\cdot\|)$. The set $\tilde{K}$ we shall call the canonical norming set of $T[(\mathcal{A}_{n_j},m_j^{-1})_{j\in\N}]$. 
\end{definition}

It is well-known that under the above assumptions the space $T[(\mathcal{A}_{n_j},m_j^{-1})_{j\in\N}]$ is a reflexive space with an unconditional basis formed by the unit vectors $(e_i)_{i=1}^\infty$ \cite[Thm I.10]{at}. Moreover, the set $\tilde{K}$ is symmetric (i.e. $f\in \tilde K$ iff $-f\in\tilde K$)  and closed under restrictions to arbitrary subsets (i.e. for any $f\in\tilde K$ and an interval $E\subset\N$ also $Ef\in \tilde K$). 

\begin{remark}\label{cite-at}
For the future reference we note that in the terminology of \cite[Section II.1]{at}, the set $\tilde K$ is equal to $W'_\kappa[G]$, with $\kappa=\omega$ and a ground set $G=\{\pm e^*_i: i\in\N\}$, and $Y_G=c_0$ with $\|\cdot\|_G=\|\cdot\|_\infty$. 
\end{remark}

Any functional $f\in\tilde{K}$ produced as in (2) above, i.e. of the form $f=m_j^{-1}(f_1+\dots+f_d)$ is called weighted, with the weight defined as $\w(f)=m_j^{-1}$. We note that the weight of a functional in $\tilde{K}$ is not necessarily uniquely determined. 

We present next the standard notion of a tree-analysis of an element of $\tilde K$. Recall that a tree  is a non-empty partially ordered  set $(\mt, \preceq)$ for which the set $\{ y \in \mt:y \preceq x \}$ is linearly ordered and finite for each $x \in \mt$. The {root} is the smallest element of the tree, terminal  nodes are the maximal elements. Given any non-terminal $t\in\mt$ by $\suc(t)$ we denote the set of immediate successors  of $t$, i.e. all nodes $r\succ t$ with no $s \in \mt$ satisfying $r \succ s \succ t$.

\begin{definition}[The tree-analysis] Let $f\in \tilde K$. By a tree-analysis of
$f$ we mean a finite family $(f_t)_{t\in \mt}\subset \tilde K$ indexed by a finite tree $\mt$ with a
unique root $0\in \mt$ such that 
\begin{enumerate}[(i)]
\item $f_{0}=f$,
\item $t\in \mt$ is terminal iff  $f_t\in \{\pm e_i^*:i\in\N\}$,
\item for every non-terminal $t\in \mt$,  $f_t=\pm m_j^{-1}E\sum_{r\in \suc(t)}f_r$, for some $j\in\N$ with $\# \suc(t)\leq n_j$, an interval $E\subset \N$, and a sequence $(f_r)_{r\in \suc(t)}\subset \tilde K$ which is block with respect to an appropriate linear ordering of $\suc(t)$.
\end{enumerate}
\end{definition}
Note that by the construction any element of $\tilde K$ admits a tree-analysis in $\tilde K$.

\subsection{Mappings on $c_{00}$} We introduce here the mappings that will be used in the definition of a space, namely the spread $S$, its adjoint $R$ and "restricted" inverse map $\Lambda$ of $R$.

We let $Y_0=\mathrm{span}\{e_{2i}:i\in\N\}\subset c_{00}$ and $S:c_{00}\to Y_0\subset  c_{00}$ be the spread $S_{\N,2\N}$, i.e. a linear bijection onto its image of the form
\[  S: c_{00}\ni \sum_{i=1}^\infty a_ie_i\mapsto \sum_{i=1}^\infty a_ie_{2i}\in Y_0\subset c_{00}  \]

We shall use also the linear adjoint of $S$ restricted to $c_{00}$, $R: c_{00}\ni f\mapsto f\circ S\in c_{00}$, with  $c_{00}$ treated as a vector spaces of functionals, i.e.  of the form
\[R: c_{00}\ni \sum_{i=1}^\infty b_ie_i^*\mapsto \sum_{i=1}^\infty b_{2i}e_i^*\in c_{00}\]
In the notation of \cite{gm2}, $R=S_{2\N,\N}$. 
\begin{remark}  \label{R properties} 
\begin{enumerate}
    \item $S$ and $R$ preserve block sequences, i.e. for any $f_1<f_2$ we have $Sf_1<Sf_2$ and $Rf_1<Rf_2$.
    \item For any $f,g\in c_{00}$ with $Rg=f$ also $R(2\range(f)g)=f$ and $2\supp(f)\subset\supp(g)$. 
    \item If $\min\range(g),\max\range(g)\in 2\N$ for $g\in c_{00}$, then $\range(g)=2\range(Rg)$.
    \item For any interval $E\subset \N$ and $f\in c_{00}$, $R(Ef)=\tilde ER(f)$, where $\tilde E=R(\chi_E)$, with $\chi_E$ - the characteristic function of $E$.     
\end{enumerate}
\end{remark}

We define a "restricted" inverse mapping to $R$, denoted by $\Lambda: c_{00}\to 2^{c_{00}}$ as
\[\Lambda(f)=\big\{(2\range(f))g: g\in R^{-1}(f)\big\}, \ \ \ f\in c_{00}.\] 
The restriction on ranges ensures that $\Lambda$ preserves block sequences, see Lemma \ref{lambda properties}\ref{lambda preserves block}.

We define inductively mappings $\Lambda^k:c_{00}\to 2^{c_{00}}$, $k\in\N_0$, as $\Lambda^0(f)=\{f\}$, $f\in c_{00}$, and 
\[\Lambda^{k+1}(f)=\bigcup_{h\in \Lambda^k(f)}\Lambda(h), \ \ f\in c_{00}, \ \  k\in\N_0\]

\begin{lemma}\label{lambda properties}
\begin{enumerate}[(1)]
\item\label{lambda equivalent}  For any $f\in c_{00}$ and $k\in\N$,  \[\Lambda^k(f)=\big\{g\in c_{00}: R^kg=f, \range(g)=2^k\range(f)\big\}\neq \emptyset\]
\item \label{lambda with R}
For any $f\in c_{00}$, $k\in\N$ and $l=1,\dots,k$, we have 
$R^l(\Lambda^k(f))=\Lambda^{k-l}(f)$.

\item\label{lambda preserves block} $\Lambda$ preserves block sequences, i.e. for any $f_1<f_2$, $g_1\in\Lambda(f_1)$, $g_2\in\Lambda(f_2)$, we have $g_1<g_2$.   
\end{enumerate}
\end{lemma}

\begin{proof} We prove  \ref{lambda equivalent} by induction on $k\in\N$. For $k=1$ the statement follow by definition of $\Lambda$ and $R$ (see Remark \ref{R properties}). Assume now that the statement holds for $k\in\N$. For any $f\in c_{00}$ we have by the definition and the inductive assumption for $k$ and for 1
\begin{align*}
\Lambda^{k+1}(f)&=\bigcup\{\Lambda(h): h\in c_{00}, R^kh=f, \range(h)=2^k\range(f)\}
\\
&=\{g\in c_{00}: Rg=h, \range(g)=2\range(h) \text{ for some }h\in c_{00} \text{ with }R^kh=f, \range(h)=2^k\range(f)\}\\
&=\{g\in c_{00}: R^{k+1}g=f, \range(g)=2^{k+1}\range(f)\}
\end{align*}
which ends the proof. 

For \ref{lambda with R} note that by \ref{lambda equivalent} for 1,  $R(\Lambda^k(f))=R(\Lambda(\Lambda^{k-1}(f)))=\Lambda^{k-1}(f)$ and proceed by induction on $l$ to obtain the statement. For \ref{lambda preserves block}  use  \ref{lambda equivalent}. 
\end{proof}

\subsection{The definition of the space $X$}

We shall use the Gowers-Maurey technique, introduced in \cite{gm} and extended in \cite{gm2} within the framework of \cite{at}. 
\begin{assumption}\label{assumption-integers}
Fix two sequences $(n_j)_j, (m_j)_j\subset 2\N$ satisfying the following: $m_1=2$, $n_1= 4$, $m_{j+1}=m_j^5$, $n_{j+1}= (5n_j)^{s_j}$, where $s_j=\log_2(m_{j+1}^3)$, for all $j\in\N$. 
\end{assumption}
Denote by $c_{00}^\mathbb{Q}$ the $\mathbb{Q}$-vector subspace of $c_{00}$ spanned by $\{e_i^*: i\in\N\}$ and fix  an injective function $\sigma: \{(f_1,\dots,f_d)\subset c_{00}^\mathbb{Q}: f_1<\dots<f_d\}\to 4\N$ such that 
$\sigma((f_1,\dots,f_d))\geq 4\max\supp (f_d)\|f_1+\dots+f_d\|_\infty^{-1}$ for any non-zero $(f_1,\dots,f_d)$.

We define now special sequences of elements of $\tilde K$ as follows. A block sequence $f_1<\dots<f_{2d}$ in $\tilde K$, $d\leq n_{2j-1}/2$, $j\in \N$, is called $j$-special, if
\begin{enumerate}[(S1)]
    \item $\w(f_1)=m_{4l-2}^{-1}$ for some $l\in\N$ with $m_{4l-2}>9n_{2j-1}^2$,
    \item $\w(f_{2i+1})=m^{-1}_{\sigma((f_1,f_2,\dots,f_{2i-1},f_{2i}))}$ for any $i=1,\dots, d-1$,
    \item $f_{2i}\in\Lambda^{k_i}(f_{2i-1})$ for some $k_i\in\N$ and $\w(f_{2i})=\w(f_{2i-1})$, for all $i=1,\dots,d$.
\end{enumerate} 

A block sequence $g_1<\dots<g_{2d}$ in $\tilde K$, with $d\leq n_{2j-1}/2$, $j\in\N$,  $k\in\N_0$, is called a $\Lambda$-$j$-special sequence $k$-modeled on a $j$-special sequence $f_1<\dots<f_{2d}$,  provided
\begin{enumerate}[($\Lambda$1)]
    \item $\w(g_i)=\w(f_i)$, for all $i=1,\dots,2d$,
    \item $g_{2i-1}\in \Lambda^k(f_{2i-1})$ for all $i=1,\dots, d$,
    \item $g_{2i}\in\Lambda^k(f_{2i})\cap \Lambda^{k_i}(g_{2i-1})$, where $f_{2i}\in\Lambda^{k_i}(f_{2i-1})$, for all $i=1,\dots,d$.
\end{enumerate}
By definition, $j$-special sequences are $\Lambda$-$j$-special sequences. 
\begin{remark}\label{special sequences properties}
\begin{enumerate}
\item Recall that the weight of a functional in $\tilde K$ is not uniquely defined. However, when we talk about elements of a $\Lambda$-special sequence, weights will refer to those defined in (S1), (S2) or ($\Lambda$1).
\item  Note the tree-like property of the family of special sequences: for any $\Lambda$-special sequences $f_1<\dots<f_{2d}$ and $h_1<\dots<h_{2a}$ and $r:=\min\{i=1,\dots,\min\{d,a\}, f_{2i-1}\neq h_{2i-1}\}$ we have $\w(f_i)\neq \w(h_s)$ for any $(s,i)\in(\{2r+1,\dots,2a\}\times\{1,\dots,2d\})\cup (\{2r-1,2r\}\times \{2r+1,\dots,2d\})$. 
\end{enumerate}
\end{remark}

Now we are ready to define an increasing sequence of sets $K_n\subset c_{00}$, $n\in\N_0$, whose union will form the norming set $K$ of the space $X$. We define sets $K_n\subset \tilde K$, $n\in\N_0$, inductively, verifying at each step the following properties.
\begin{enumerate}[(K1)]
\item $K_n$ is symmetric and closed under restriction to intervals. 
\item Any $f\in K_n$ has a tree-analysis $(f_t)_{t\in\mt}\subset K_n$. 
\item $R(K_n)=K_n$.
\end{enumerate}
Let $K_0=\{\pm e^*_i: i\in \N\}$. Conditions (K1)-(K3) are satisfied. 

Fix $n\in\N$ and assume that we have defined $K_{n-1}$, satisfying in particular conditions (K1)-(K3). 
Let $K_n$ be the collection of all functionals $f$ of one of the forms listed below. 
\begin{enumerate}
\item $f\in K_{n-1}$. 
\item $f=m_{2j}^{-1}(f_1+\dots+f_d)$, with $j\in\N$ and a block sequence $f_1<\dots<f_d$, $d\leq n_{2j}$, of elements of  $K_{n-1}$. The resulting  functional is called regular. 
\item $f=\pm Em_{2j-1}^{-1}R^k(f_1+\dots+f_{2d})$, with  $k\in\N_0$, an interval $E\subset\N$ and a $j$-special sequence  $f_1<\dots<f_{2d}$ of elements of $K_{n-1}$. The resulting functional is called $R$-special. 
\item $f=\pm Em_{2j-1}^{-1}(g_1+\dots+g_{2d})$, for an interval $E\subset\N$ and a $\Lambda$-$j$-special sequence $g_1<\dots<g_{2d}$ of elements of $K_{n-1}$ modeled on some $j$-special sequence of elements of $K_{n-1}$. The resulting functional is called $\Lambda$-special. 
    
\end{enumerate}
Both $R$-special and $\Lambda$-special functionals will be called special. Note that a functional of the form $\pm Em_{2j-1}(f_1+\dots+f_{2d})$, for some $j$-special sequence $f_1<\dots<f_{2d}$ is both $R$-special and $\Lambda$-special. 

We verify now conditions (K1)-(K3) for $K_n$. 

(K1) First note that including operations "sign-change" and "restriction to an interval" is redundant in (2), as by the inductive assumption and definition of $K_n$, for any regular $f$ and an interval $E\subset\N$,  $-f$ and $Ef$ are regular functionals also. Therefore (K1) follows by the definition of $K_n$ and the inductive assumption.  

(K2) follows immediately by the definition of $K_n$, the inductive hypothesis and properties (K2) and (K3) of $K_{n-1}$. 

(K3) We show first the inclusion $R(K_n)\subset K_n$ in (K3). For functionals from $K_{n-1}$, regular or $R$-special use (K1),(K3) for $K_{n-1}$ and Remark \ref{R properties}(1),(4). Assume now that $f$ is a $\Lambda$-special functional generated by a $\Lambda$-special sequence $g_1<\dots<g_{2d}$ $k$-modeled on some special sequence for some $k\in\N$ (the case $k=0$ is covered by the case of $R$-special functionals). We claim that 
$Rg_1<\dots<Rg_{2d}$ is a $\Lambda$-special sequence of elements of $K_{n-1}$, $(k-1)$-modeled on the same special sequence, which by linearity of $R$ and Remark \ref{R properties}(1),(4) proves that $f\in R(K_n)$. Note first that  $Rg_i\in K_{n-1}$ by (K3) for $K_{n-1}$, $\w(Rg_i)=\w(g_i)=w(f_i)$, and $Rg_i\in\Lambda^{k-1}(f_i)$, $i=1,\dots, 2d$, by Lemma \ref{lambda properties} \ref{lambda with R}. In order to prove $Rg_{2i}\in\Lambda^{k_i}(Rg_{2i-1})$, where $g_{2i}\in\Lambda^{k_i}(g_{2i-1})$,  $i=1,\dots, d$, we shall use Lemma \ref{lambda properties} \ref{lambda equivalent}. First note that $R^{k_i}(Rg_{2i-1})=R(R^{k_i}g_{2i}) =Rg_{2i}$ by Lemma \ref{lambda properties}(1). Moreover, as $k\in\N$, $\min\range(g_{2i-1}),\max\range(g_{2i-1}), \min\range(g_{2i}), \max\range(g_{2i})\in 2\N$, thus by Remark \ref{R properties}(3) and Lemma \ref{lambda properties}(1),  $\range(Rg_{2i})=2\range(g_{2i})=2^{k_i+1}\range(g_{2i-1})=2^{k_i}\range(Rg_{2i-1})$. Therefore, once again applied Lemma \ref{lambda properties} \ref{lambda equivalent} ends the proof. 

We proceed to the reverse inclusion, i.e. $K_n\subset R(K_n)$ in (K3). Fix $f\in K_n$. We shall prove that $R^{-1}(f)\cap K_n\neq\emptyset$. Assume $f$ is regular, i.e. $f=m_{2j}^{-1}(f_1+\dots+f_d)$, with $f_1<\dots<f_d$ in $K_{n-1}$. By (K3) and (K1) for $K_{n-1}$ pick $h_i\in \Lambda(f_i)\cap K_{n-1}$, $i=1,\dots,d$. By Lemma \ref{lambda properties} \ref{lambda preserves block} $h:=m_{2j}^{-1}(h_1+\dots+h_d)\in K_n$, whereas $Rh=f$ (use linearity of $R$ and Lemma \ref{lambda properties} \ref{lambda equivalent}). 

Any $R$-special functional with $k\in\N$ (in the notation of (3) in the definition of $K_n$) belongs to $R(K_n)$ by definition of $K_n$ and Remark \ref{R properties}(4). 

Finally assume $f$ is $\Lambda$-special, i.e. $f=\pm Em_{2j-1}^{-1}(g_1+\dots+g_{2d})$, for a $\Lambda$-special sequence  $g_1<\dots<g_{2d}$ in $K_{n-1}$, $k$-modeled on a special sequence $f_1<\dots<f_{2d}$ of elements of $K_{n-1}$ for some $k\in\N_0$. We shall need the following observation. 

\begin{remark}\label{remark claim}
For any regular $\tilde g\in K_{n-1}$ there is $\tilde h\in \Lambda(\tilde g)\cap K_{n-1}$ with $\w(\tilde h)=\w(\tilde g)$.  
Indeed, let $\tilde g=m_{2l}^{-1}(\tilde g_1+\dots+\tilde g_s)$ for some $\tilde g_1<\dots<\tilde g_s\in K_{n-2}$. Pick $\tilde h_r\in \Lambda(\tilde g_r)\cap K_{n-2}$ by (K3) and (K1) for $K_{n-2}$, $r=1,\dots,s$. Then by Lemma \ref{lambda properties} \ref{lambda equivalent} $\tilde h:=m_{2l}^{-1}(\tilde h_1+\dots+\tilde h_s)\in \Lambda(\tilde g)\cap K_{n-1}$. 

Note that we use in the justification only properties (K1), (K3) and the operation described in (2) of the definition of $K_{n-1}$ producing regular functionals.
\end{remark}

Coming back to the proof of (K3) for $K_n$, let $k_i\in\N$ be such that $g_{2i}\in\Lambda^{k_i}(g_{2i-1})$, $i=1,\dots,d$ (by ($\Lambda$3)). By Remark \ref{remark claim} pick $h_{2i}\in \Lambda(g_{2i})\cap K_{n-1}$ with $\w(h_{2i})=\w(g_{2i})$, and let $h_{2i-1}=R^{k_i-1}g_{2i}\in K_{n-1}$ (by (K3) for $K_{n-1}$), for each $i=1,\dots,d$. Note that $f=Rh$ for $h:=\pm(2E)(h_1+\dots+h_{2d})$ by the linearity of $R$, Lemma \ref{lambda properties} \ref{lambda equivalent}) and Remark \ref{R properties}(4). We claim that $h_1,\dots,h_{2d}$ is a $\Lambda$-special sequence $(k+1)$-modeled on the same special sequence $f_1<\dots<f_{2d}$, which will finish the proof. First note that $h_{2i}\in\Lambda(g_{2i})$ by definition and $h_{2i-1}\in R^{k_i-1}(\Lambda^{k_i}(g_{2i-1}))=\Lambda(g_{2i-1})$ by Lemma \ref{lambda properties} \ref{lambda with R}, $i=1,\dots,d$, thus by Lemma \ref{lambda properties} \ref{lambda preserves block} $h_1<\dots<h_{2d}$. It follows also that $h_i\in\Lambda(g_i)\subset \Lambda^{k+1}(f_i)$, $i=1,\dots, 2d$, as $g_1<\dots<g_{2d}$ is $k$-modeled on $f_1<\dots<f_{2d}$. By the choice of $h_i$'s we have $\w(h_i)=\w(g_i)=\w(f_i)$ for each $i=1,\dots, 2d$. It remains to show  that $h_{2i}\in\Lambda^{k_i}(h_{2i-1})$, $i=1,\dots, d$, we shall use Lemma \ref{lambda properties} \ref{lambda equivalent}. By the definition and Lemma \ref{lambda properties}(1), $h_{2i-1}=R^{k_i-1}g_{2i}=R^{k_i}h_{2i}$. As $h_{2i}\in\Lambda(g_{2i})\subset\Lambda^{k_i+1}(g_{2i-1})$, $\range(h_{2i})=2^{k_i+1}\range(g_{2i-1})$ by Lemma \ref{lambda properties}(1). On the other hand $h_{2i-1}\in\Lambda(g_{2i-1})$, as we proved above, thus $\range(h_{2i-1})=2\range(g_{2i-1})$. Therefore $\range(h_{2i})=2^{k_i}\range(h_{2i-1})$ and Lemma \ref{lambda properties} \ref{lambda equivalent} ends the proof and thus the whole inductive construction.

\

Let $K=\bigcup_{n=0}^\infty K_n\subset\tilde K$. Let $\|\cdot\|$ be the norm on $c_{00}$ with the norming set $K$, i.e. $\|\cdot\|=\sup\{|f(\cdot)|:f\in K\}$. Let $X$ be the completion of $(c_{00}, \|\cdot\|)$ and $Y=\overline{Y_0}\subset X$. 

By the  definition of $K$ and $X$ we have the following. 
\begin{fact}\label{fact after def X}
 \begin{enumerate}
\item As $K_0\subset K$, $\|\cdot\|\leq \|\cdot\|_\infty$. 
\item $K$ satisfies properties (K1)-(K3).
\item By (K1), the unit vectors form a bimonotone basis of $X$. 
\item For any $j\in\N$ and a block sequence  $f_1<\dots<f_d$, $d\leq n_{2j}$, of elements of $K$, also $m_{2j}^{-1}(f_1+\dots+f_d)\in K$. 
 \end{enumerate}   
\end{fact}

\begin{remark}\label{remark after def X}
We recall that in the construction of \cite{gm2} the norming set was closed both under $R$ and $S$ applied to functionals, which generate also bounded projection on $Y$ in the resulting space. In order to avoid it, we close the norming set under $R$ and some selections of  $\Lambda$, typically different from $S$ (cf. the definition of $\Lambda$-special sequences). This precise choice of selections of $\Lambda$ allowed into $K$ permits an extensive use of Gowers-Maurey technique that guarantees the subspace $Y$ is uncomplemented in $X$ (see Section 3.). 
\end{remark}

\begin{proposition}\label{isometric}
The injection $S:c_{00}\to Y_0\subset c_{00}$ is an isometry with the norm $\|\cdot\|$, thus extends to an isometry $X\to X$ with $S(X)=Y$.
\end{proposition}
\begin{proof}
Fix $x\in c_{00}$. 
Given $f\in K$, by (K3) pick $g\in  K$ with $f=Rg$ and estimate $|f(x)|=|(Rg)(x)|=|g(Sx)|\leq \|Sx\|$. On the other hand, for any $g\in K$ also $Rg\in K$ by the property (K3), thus $|g(Sx)|=|(Rg)(x)|\leq \|x\|$. Thus we obtain $\|x\|=\|Sx\|$ for any $x\in c_{00}$, which ends the proof. 
\end{proof}

\section{Classical Gowers-Maurey tools}

We recall here the notions and results from \cite{at}. First note, that $K$ ($=D_G$ in the notation of \cite{at}) by properties (K1), (K2) and (4) of Fact \ref{fact after def X}, is a tree-complete extension of the ground set $G:=\{\pm e^*_i: i\in\N\}$ (see \cite[Def. II.10]{at}), therefore we can directly cite results from \cite{at}. The space $X$ defined in the previous section is equal to the space $\mathfrak{X}_\omega(D_G)$  of \cite[Section II.2]{at}.

Recall that a bounded operator $A:U\to V$ between Banach spaces $U,V$ is strictly singular, if none of its restrictions to an infinite dimensional subspaces of $U$ is an isomorphism onto its image. Equivalently, for any infinite dimensional subspace $Z$ of $U$ and any $\varepsilon>0$ there is a vector $z\in Z$ with $\|z\|>1$ and $\|Az\|<\varepsilon$. 
\begin{proposition}\label{cite-at-ss-extension}
The identity operator $X\to c_0$ is strictly singular. 
\end{proposition}
\begin{proof}
By the Krein-Milman-Rutman Theorem we can consider only block subspaces in $X$. In a given block subspace, for $j\in\N$, pick a normalized block sequence $x_1<\dots<x_{n_{2j}}$ and a block sequence  $f_1<\dots<f_{n_{2j}}$ of elements of $K$ with $f_i(x_i)=1$ for any $i=1,\dots, n_{2j}$. Let $x=m_{2j}n_{2j}^{-1}(x_1+\dots+x_{n_{2j}})$. Then, by the definition of $K$, also $f=m_{2j}^{-1}(f_1+\dots+f_{n_{2j}})\in K$, thus $\|x\|\geq f(x)=1$, whereas $\|x\|_\infty=m_{2j}n_{2j}^{-1}\max\{\|x_i\|_\infty: i=1,\dots,n_{2j}\}\leq \max\{\|x_i\|, i=1,\dots,n_{2j}\}= m_{2j}n_{2j}^{-1}$ (see Fact \ref{fact after def X}). By Assumption \ref{assumption-integers},  $m_{2j}n_{2j}^{-1}\to 0$, as $j\to\infty$, which ends the proof.      
\end{proof}
The above observation stated in the terminology of \cite[Section II.2]{at} claims that $X$ is a strictly singular extension of $Y_G(=c_0)$ (cf. Remark \ref{cite-at}).

\begin{definition}[$\ell_1$-averages]\cite[Def. II.21]{at}
A vector $x\in X$  is called a $C$-$\ell_1^N$-average, $C>1$, if $\|x\|=1$ and there is a block sequence $(x_n)_{n=1}^N\subset X$ with $x=N^{-1}(x_1+\dots+x_N)$ and $\|x_n\|\leq C$ for each $n=1,\dots,N$.  
\end{definition} 
If $C=2$ we will omit it in the notation and say  that $x$ is an $\ell_1^N$-average. 
By Corollary \ref{cite-at-ss-extension} we have the following 
\begin{lemma}\label{av-existence}\cite[Lemma II.22]{at}  Any block subspace of $X$ for any $N\in\N$ and $\varepsilon>0$ contains an $\ell_1^N$-average $x$ with $\|x\|_\infty<\varepsilon$. 
\end{lemma}

\begin{lemma}\label{av-est}\cite[Lemma II.23]{at}
Let $x\in X$ be an $\ell_1^{n_j}$-average, for some $j\in\N$. Then for any $f\in K$ with $\w(f)>m_j^{-1}$ we have
$|f(x)|\leq 3\w(f)$.
\end{lemma}

\begin{definition}[RIS]\cite[Def. II.13]{at}\label{ris-def} A block sequence $(x_i)_{i\in I}\subset
X$ is called a $(C,\varepsilon)$- {rapidly increasing sequence} (RIS), for $C>1$, $\varepsilon>0$ and an interval $I\subset\N$, if there is a strictly increasing
sequence $(j_i)_{i\in I}\subset L_0$ such that 

\begin{enumerate}

\item $\|x_i\|\leq C$, $\|x_i\|_\infty<\varepsilon$, for any $i\in I$,

\item 
$2m_{j_{i+1}}^{-1}\#\supp(x_i)<\varepsilon$ for any $i\in I\setminus\{\max I\}$,

\item $| f(x_i)|\leq C\w(f)$ for any weighted $f\in K$ with $\w(f)> m_{j_i}^{-1}$ for any $i\in I$.

\end{enumerate}
\end{definition}

Following the proof of \cite[Prop. II.25]{at} we obtain the following.
\begin{lemma}\label{ris-existence}
Any block subspace of $X$ contains for any $\varepsilon>0$ a normalized block sequence $(x_i)_{i=1}^\infty$ 
such that   $(S^kx_i)_{i=1}^\infty$ is a $(3,\varepsilon)$-RIS for any $k\in\N_0$.
\end{lemma}
\begin{proof} 
In a given block subspace of $X$  pick inductively (using Lemma \ref{av-existence}) a block sequence $(x_i)_{i=1}^\infty$ of $\ell_1$-averages, so that for some  $(j_i)_{i=1}^\infty\subset \N$ we have the following:
\begin{enumerate}[(i)]
    \item $x_i$ is an $\ell_1^{j_i}$-average with $\|x_i\|_\infty<\varepsilon$, for all $i\in\N$,
    \item $(x_i)_{i=1}^\infty$ and $(j_i)_{i=1}^\infty$ satisfy the condition (2) of Def. \ref{ris-def}.
\end{enumerate}  
Note that for any $k\in\N$,  $(S^kx_i)_{i=1}^\infty$ is also a block sequence of $\ell_1^{j_i}$-averages (as $S$ is an isometry), that satisfies (i) and (ii). By Lemma \ref{av-est} each sequence $(S^kx_i)_{i=1}^\infty$, $k\in\N_0$, satisfies the condition (3) of Def. \ref{ris-def}, thus is a $(3,\varepsilon)$-RIS. 
\end{proof}

\begin{proposition}\cite[Prop. II.19]{at}\label{basic-inequality-cor} Fix $j\in\N$. Let $(x_i)_{i=1}^d\subset X$, $d\leq n_{2j-1}$, be a $(C,m_j^{-2})$-RIS. Then for any weighted functional $f\in K$  we have
$$
\Big| f\Big(\frac{1}{n_j}\sum_{i=1}^dx_i\Big)\Big|\leq \begin{cases} \frac{2C}{m_j^2} \ \ \ \ \ \ \ \ \text{if}\ \ \ \w(f)<\frac{1}{m_j} \\
\frac{2C}{m_j} \ \ \ \ \ \ \ \ \text{if}\ \ \ \w(f)=\frac{1}{m_j}\\
\frac{3C}{m_j}\w(f) \ \ \ \text{if}\ \ \ \w(f)>\frac{1}{m_j}
\end{cases}
$$
In particular
$$
\Big\|\frac{1}{n_j}\sum_{i=1}^dx_i\Big\|\leq \frac{2C}{m_j}
$$
Moreover, if for any weighted functional $f\in K$ with $\w(h)=m_j^{-1}$ and any interval $E\subset \{1,\dots,d\}$,
\[\Big|h\Big(\sum_{i\in E}x_i\Big)\Big|\leq C, \]
then
\[\Big\|\frac{1}{n_j}\sum_{i=1}^dx_i\Big\|\leq \frac{4C}{m_j^2}\]
\end{proposition}
\begin{corollary}
\label{reflexive}\cite[Prop. II.28]{at} The basis $(e_i)_{i=1}^\infty$ of $X$ is boundedly complete and shrinking, thus the space $X$ is reflexive.   
\end{corollary}

\section{The subspace $Y$ is  uncomplemented in $X$}

This section contains the proof of uncomplementability of $Y$ in $X$, using the  classical Gowers-Maurey technique within the framework of \cite{at}. More complicated forms of special sequences used in the definition of $X$ require additional technical assumptions on so-called dependent sequences (see Definition \ref{def dep}) and more technical reasoning we present in detail below.

\begin{definition}[exact pairs]\cite[Def. II.31]{at}\label{def-exact}
A pair $(x,f)$, with  $x\in X$, $f\in K$, is called a $j$-exact pair, $j\in \N$, provided 
\begin{enumerate}
\item $f$ is weighted with $\w(f)=m_j^{-1}$,
\item $1\leq \|x\|\leq 6$, $\|x\|_\infty<m_j^{-1}$,
\item $f(x)=1$,  $\range(f)=\range(x)$,
\item $|h(x)|\leq 9\max\{m_j^{-1},\w(h)\}$ for any weighted $h\in K$ with weight $\w(h)\neq m_j^{-1}$. 
\end{enumerate}
A pair $(x,f)$, with $x\in X$, $f\in K$, is a called a stable $j$-exact pair, $j\in\N$, if $(S^kx,g)$ is a $j$-exact pair for any $k\in\N_0$ and $g\in \Lambda^k(f)\cap K$ with $\w(g)=\w(f)$.

\end{definition}
Following the proof of \cite[Prop. II.32]{at} we obtain the following. 
\begin{lemma}\label{exact existence}
Any block subspace of $X$, for any $j\in\N$ and $\varepsilon>0$, contains a vector $x$ such that for some $f\in K$, $(x,f)$ is a stable $2j$-exact pair.
\end{lemma}
\begin{proof} In a given block subspace of $X$ pick (by Lemma \ref{ris-existence}) a normalized block sequence $(x_i)_{i=1}^{n_{2j}}$ with   $(S^kx_i)_{i=1}^{n_{2j}}$ a $(3,m_{2j}^{-1})$-RIS for any $k\in\N_0$. 

For each $i=1,\dots,n_j$ pick $f_i\in K$ with $f_i(x_i)=1$ and $\range(f_i)=\range (x_i)$ (as the basis in $X$ is bimonotone). Let 
\[x=\frac{m_{2j}}{n_{2j}}\sum_{i=1}^{n_{2j}}x_i, \ \ \ f=\frac1{m_{2j}}\sum_{i=1}^{n_{2j}}f_i\in K\]

By Prop. \ref{basic-inequality-cor}, as $S$ is an isometry, $1\leq\|S^kx\|\leq 6$ for all $k\in\N_0$. Moreover, $\|S^kx\|_\infty<m_{2j}^{-1}$ for all $k\in\N_0$. 

Note that for  any $k\in\N$ and  $g\in\Lambda^k(f)\cap K$, $g_k(S^kx)=(R^kg)(x)=f(x)=1$ and $\range(g)=2^k\range(f)=2^k\range(x)=\range(S^kx)$ by Lemma \ref{lambda properties} \ref{lambda equivalent}. 

We proceed to the proof of (4) of Def. \ref{def-exact}. Note first that $x$ satisfies (4) by Prop. \ref{basic-inequality-cor}. Now for any $k\in\N$ and $h\in K$, by the definition of $R$,  $h(S^kx)=(R^kh)x$. As $R$ preserves the weight of a functional, vectors $S^kx$, $k\in\N$, also satisfy the condition (4), which ends the proof. 
\end{proof}

\begin{definition}[Dependent sequences]\label{def dep} A double sequence $(x_i,f_i)_{i=1}^{n_{2j-1}}\subset
X\times K$ is called a $j$-dependent sequence, $j\in\N$, if for some sequences $(j_i)_{i=1}^{n_{2j-1}/2}\subset 2\N$ and $(k_i)_{i=1}^{n_{2j-1}/2}\subset \N_0$ we have the following
\begin{enumerate}[(D1)]
\item\label{dep-seq-vectors} 
$m_{j_1}>9n_{2j-1}^2$,
\item\label{dep-seq-special} $(f_i)_{i=1}^{n_{2j-1}}$ is a $j$-special sequence, with $f_{2i}\in\Lambda^{k_i}(f_{2i-1})$ and $\w(f_{2i-1})=m_{j_i}^{-1}$ for any $i=1,\dots,n_{2j-1}/2$, 
\item\label{dep-seq-vectors-shifted} $x_{2i}=S^{k_i}(x_{2i-1})$, for any $i=1,\dots,n_{2j-1}/2$,
\item\label{dep-seq-exact} $(x_{2i-1},f_{2i-1})$ and $(x_{2i}, f_{2i})$ are $j_i$-exact pairs, for any $i=1,\dots,n_{2j-1}/2$,
\item\label{dep-seq-limit-intersection} for any $k\in\N_0$ there is at most one $i\in\{1,\dots,n_{2j-1}/2\}$ with $h(x_{2i})\neq 0$ for some $h\in\Lambda^k(f_{2i-1})$ and at most one $s\in\{1,\dots,n_{2j-1}/2\}$ with $R^kf_{2s}(x_{2s-1})\neq 0$. 
\end{enumerate}
\end{definition}

\begin{remark}\label{remark dep}
Note that by (D1), (D2), (D4) and the choice of $\sigma$, we have 
\begin{enumerate}
        \item[(D6)] $4m_{j_{i+1}}^{-1}\#\supp(x_{2i-1})<m_{2j-1}^{-2}$   for any $i=1, \dots, n_{2j-1}-1$. 
    \end{enumerate}
\end{remark}

\begin{lemma}\label{lemma dep} Let $(x_i)_{i=1}^{n_{2j-1}}\subset c_{00}$, $f_i)_{i=1}^{n_{2j-1}}\subset K$ be block sequences with $\range(f_{2i-1})=\range(x_{2i-1})$, $f_{2i}\in\Lambda^{k_i}(f_{2i-1})$, $x_{2i}=S^{k_i}(x_{2i-1})$, $i=1,\dots,n_{2j-1}/2$ for some $(k_i)_{i=1}^{n_{2j-1}/2}\subset \N$. Assume that for some $(s_i)_{i=1}^{n_{2j-1}/2}\subset\N$,
\begin{enumerate}
    \item[(D7)]   $\range(x_{2i-1})\subset [2, 2^{s_i}]$ and $k_i+s_i<k_{i+1}-s_{i+1}$ for any $i=1,\dots,n_{2j-1}$. 
\end{enumerate}
Then $(x_i,f_i)_{i=1}^{n_{2j-1}}$ satisfies  (D5).
\end{lemma}
\begin{proof}
Assume (D7) and fix $k\in\N$. With the above notation for any  $i=1,\dots,n_{2j-1}$ we have $\range(f_{2i})=\range(x_{2i})\subset [2^{k_i}, 2^{k_i+s_i}]$,  $\range(h)\subset [2^{k},2^{k+s_i}]$ for any $h\in\Lambda^k(f_{2i-1})$ by Lemma \ref{lambda properties} \ref{lambda equivalent} and $\range(R^kf_{2i})\subset[\max\{1,2^{k_i-k}\}, \max\{1,2^{k_i+s_i-k}\}]$. 

Thus if $h(x_{2i})\neq 0$ for some $h\in\Lambda^k(f_{2i-1})$, then $k\leq k_i+s_i$ and $k_i\leq k+s_i$, i.e. $k\in [k_i-s_i, k_i+s_i]$. By (D7) there is at most one $i$ with such property, thus with $h(x_{2i})\neq 0$ for some $h\in\Lambda^k(f_{2i-1})$. 

Also, if $(R^kf_{2i})(x_{2i-1})\neq 0$, then 
$0\leq k_i+s_i-k$ and $k_i-k\leq s_i$, i.e. $k\in  [k_i-s_i, k_i+s_i]$. Thus again by (D7) there is at most one $i$ with $(R^kf_{2i})(x_{2i-1})\neq 0$. 
\end{proof}

\begin{proposition}\label{dep-est} Let $(x_i,f_i)_{i=1}^{n_{2j-1}}$ be a $j$-dependent
sequence, $j\in\N$. Then
\[\Big\|\frac{1}{n_{2j-1}}\sum_{i=1}^{n_{2j-1}}(-1)^ix_i\Big\|\leq \frac{240}{m_{2j-1}^2}\]
\end{proposition}

\begin{proof}
Notice that the sequence $(x_{2i}-x_{2i-1})_{i=1}^{n_{2j-1}/2}$ is itself a $(18,m_{2j-1}^{-2})$-RIS by (D4) and (D6).

By Prop.  \ref{basic-inequality-cor} it is enough to show that for any special $h\in K$ with $\w(h)=m_{2j-1}^{-1}$ and any interval $E$ we have $| h(\sum_{i\in E}(x_{2i}-x_{2i-1}))|\leq 60$. Fix such $h\in K$  and an interval $E\subset\N$. 

Assume first that $h$ is $R$-special of the form 
\[h=m_{2j-1}^{-1}(FR^k(f_{2l-1}+f_{2l})+\dots+R^k(f_{2r-1}+f_{2r})+R^kh_{2r+1}+\dots+R^kh_d)\] for $k\in\N_0$, an interval $F\subset\N$ and a $j$-special sequence
$(f_1,\dots,f_{2r},h_{2r+1},\dots,h_{n_{2j-1}})$, $1\leq l\leq r$, $2r\leq
d\leq n_{2j-1}$, with $h_{2r+1}\neq
f_{2r+1}$  (for other forms the estimates are even simpler). 

Note the following estimates. 
For any $(s,i)\in (\{2r+3,\dots,d\}\times \{1,\dots,n_{2j-1}\})\cup (\{2r+1,2r+2\}\times \{2r+3,\dots,n_{2j-1}\})$, by Remark \ref{special sequences properties}(2),  $\w(R^kh_s)\neq \w(f_i)$, thus $|
(R^kh_s)(x_i)|\leq n_{2j-1}^{-2}$ by (D4) and (D6). For any $i\neq s$ by (D2) $\w(f_{2s-1})\neq m_{j_i}^{-1}$, hence $|R^k(f_{2s-1}+f_{2s})(x_{2i}-x_{2i-1})|\leq 4 n_{2j-1}^{-2}$ by (D4) and (D6). Thus, using the estimate $\|x_i\|\leq 6$ for any $i$ by (D4), we obtain 
\begin{align*}
\Big| h\Big(\sum_{i\in E}(x_{2i}-x_{2i-1}\Big)\Big| & \leq 5+|FR^k(f_{2l-1}+f_{2l})(x_{2l}-x_{2l-1})|+|R^k(h_{2r-1}+h_{2r})(x_{2r}-x_{2r-1})|\\
&+\sum_{i=l+1}^r|R^k(f_{2i-1}+f_{2i})(x_{2i}-x_{2i-1})|
\\
&\leq 53+\sum_{i=l+1}^r|(R^kf_{2i})(x_{2i})-(R^kf_{2i-1})(x_{2i-1})|+\sum_{i=1}^{l+1}|(R^kf_{2i})(x_{2i-1})|
\\
&+\sum_{i=1}^{l+1}|(R^kf_{2i-1})(x_{2i})|
\end{align*}

Notice that $(R^kf_{2i})(x_{2i})=(R^kf_{2i})(S^{k_i}x_{2i-1})=(R^{k+k_i}f_{2i})(x_{2i-1})=(R^kf_{2i-1})(x_{2i-1})$ for any $i=1,\dots,r$ by  (D2), (D3) and Lemma \ref{lambda properties}(1). 

As $\min\supp (x_{2i})>\max\supp(x_{2i-1})=\max\supp(f_{2i-1})\geq \max\supp( R^kf_{2i-1})$ for any $i$ by (D4) and the definition of $R$, $(R^kf_{2i-1})(x_{2i})=0$ for any $i=l+1,\dots,r$.

Finally, by (D5) there is at most one $s$ with $(R^kf_{2s})(x_{2s-1})\neq 0$, thus we obtain 
\[
\Big|h\Big(\sum_{i\in E}(x_{2i}-x_{2i-1})\Big)\Big|\leq 53 + 0+ 0+6=59
\]

The proof in the case of $\Lambda$-special functionals is analogous. Let $h$ be a $\Lambda$-special functional of the form 
\[h=m_{2j-1}^{-1}(F(g_{2l-1}+g_{2l})+\dots+(g_{2r-1}+g_{2r})+g_{2r+1}+\dots+g_d)\] 
for $k\in\N$, an interval $F\subset\N$ and a $\Lambda$-$j$-special sequence $g_1<\dots<g_{n_{2j-1}}$ $k$-modeled on a $j$-special sequence 
$(f_1,\dots,f_{2r},h_{2r+1},\dots,h_{n_{2j-1}})$, for $k\in\N$, $1\leq l\leq r$, $2r\leq
d\leq n_{2j-1}$, with $h_{2r+1}\neq
f_{2r+1}$  (for other forms the estimates are even simpler).  

As before, for any $(s,i)\in (\{2r+3,\dots,d\}\times \{1,\dots,n_{2j-1}\})\cup (\{2r+1,2r+2\}\times \{2r+3,\dots,n_{2j-1}\})$, by Remark \ref{special sequences properties}(2),  $\w(g_s)\neq \w(f_i)$, thus $|
g_s(x_i)|\leq n_{2j-1}^{-2}$ by (D4) and (D6). For any $i\neq s$ by (D2) $\w(g_{2s-1})\neq m_{j_i}$, thus $|(g_{2s-1}+g_{2s})(x_{2i}-x_{2i-1})|\leq 4 n_{2j-1}^{-2}$ by (D4) and (D6). Thus, as $\|x_i\|\leq 6$ for any $i$ by (D4), we have 
\begin{align*}
\Big| h\Big(\sum_{i\in E}(x_{2i}-x_{2i-1}\Big)\Big| & \leq 5+|F(g_{2l-1}+g_{2l})(x_{2l}-x_{2l-1})|+|(g_{2r-1}+g_{2r})(x_{2r}-x_{2r-1})|\\
&+\sum_{i=l+1}^r|(g_{2i-1}+g_{2i})(x_{2i}-x_{2i-1})|
\\
&\leq 53+\sum_{i=l+1}^r|g_{2i}(x_{2i})-g_{2i-1}(x_{2i-1})| +\sum_{i=1}^{l+1}|g_{2i}(x_{2i-1})|
\\
&+\sum_{i=1}^{l+1}|g_{2i-1}(x_{2i})|
\end{align*}

Notice that $g_{2i}(x_{2i})=g_{2i}(S^{k_i}x_{2i-1})=R^{k_i}g_{2i}(x_{2i-1})=g_{2i-1}(x_{2i-1})$ for any $i=1,\dots,r$ by (D2), (D3), ($\Lambda$3) and Lemma \ref{lambda properties}(1). 

As $\min\supp(g_{2i})\geq\min\supp(f_{2i})=\min\supp (x_{2i})>\max\supp(x_{2i-1})$ for any $i$ by (D4) and the definition of $\Lambda$-special sequences, $g_{2i}(x_{2i-1})=0$ for any $i=l+1,\dots,r$.

Finally, by (D5) there is at most one $i$ with $g_{2i-1}(x_{2i})\neq 0$, thus we obtain 
\[
\Big|h\Big(\sum_{i\in E}(x_{2i}-x_{2i-1})\Big)\Big|\leq 53 + 0+ 6+0=59
\]
which ends the proof. 
 \end{proof}

\begin{proposition}\label{dep existence} For any block subspace $Z$ of $X$ and $j\in\N$ there is a $j$-dependent sequence 
$(x_i,f_i)_{i=1}^{n_{2j-1}}$ with
$(x_{2i})_{i=1}^{n_{2j-1}/2}\subset Y$ and
$(x_{2i-1})_{i=1}^{n_{2j-1}/2}\subset Z$. \end{proposition}

\begin{proof} Fix $j\in\N$. 
We pick inductively on $i$, using Lemma \ref{exact existence},  block sequences $(x_i)_{i=1}^{n_{2j-1}}$, $(f_i)_{i=1}^{n_{2j-1}}$,  $(j_i)_{i=1}^{n_{2j-1}/2}\subset 2\N$, $(k_i)_{i=1}^{n_{2j-1}/2}, (s_i)_{i=1}^{n_{2j-1}/2}\subset \N$,  with $(x_{2i-1})_{i=1}^{n_{2j-1}/2}\subset Z$, $(x_{2i})_{i=1}^{n_{2j-1}/2}\subset Y$,  satisfying conditions  (D1)-(D4) and (D7).

Pick any $j_1\in 4\N-2$ with $m_{j_1}>9n_{2j-1}^2$ and, by Lemma \ref{exact existence}, a stable $j_1$-exact pair $(x_1,f_1)$ with $x_1\in Z$, $\min\supp(x_1)\geq 2$. Pick $k_1\in\N$ such that $x_2>x_1$, where $x_2=S^{k_1}x_1\in Y$. 
As in Remark \ref{remark claim} pick $f_2\in\Lambda^{k_i}(f_1)\cap K$ with $\w(f_2)=\w(f_1)$. Choose  $s_1\in\N$ with $\range(x_1)\subset [2,2^{s_1}]$. 

Assume that $(x_l)_{l=1}^{2i}$, $(f_l)_{l=1}^{2i}$, $(j_l)_{l=1}^i$, $(k_l)_{l=1}^i$ and $(s_l)_{l=1}^i$, $i<n_{2j-1}/2$, satisfying the required conditions, are chosen. 

Let $j_{i+1}=\sigma((f_1,f_2,\dots,f_{2i-1}, f_{2i}))$. Choose by Lemma \ref{exact existence} a stable $j_{i+1}$-exact pair  $(x_{2i+1},f_{2i+1})$ with $x_{2i+1}\in Z$, $x_{2i+1}>x_{2i}$. Pick $s_{i+1}\in\N$ with $\range(x_{2i-1})\subset [0,2^{s_{i+1}}]$. Pick $k_{i+1}> k_i+s_i+s_{i+1}$ with  $x_{2i+2}>x_{2i+1}$, where  $x_{2i+2}=S^{k_{i+1}}x_{2i+1}$. Again as in Remark \ref{remark claim} choose $f_{2i+2}\in \Lambda^{k_{i+1}}(f_{2i+1})\cap K$ with $\w(f_{2i+2})=\w(f_{2i+1})$.  

By construction conditions (D1)-(D4) and (D7) are satisfied, which ends the inductive construction and thus, by Lemma \ref{lemma dep}, the proof. 
\end{proof}

\begin{theorem}\label{not-complemented} The subspace $Y$ is not complemented in $X$.   
\end{theorem}

\begin{proof} We shall prove that for any infinite dimensional subspace $Z$ of $X$ and $\varepsilon>0$ there are vectors $y\in Y$, $z\in Z$ with $\|y+z\|\geq 1$ and $\|y-z\|<\varepsilon$. By the Krein-Milman-Rutman Theorem (\cite{lt}) it is enough to consider only block subspaces of $X$. 

Fix a block subspace  $Z$ of $X$ and  $j\in\N$. By Lemma \ref{dep existence} pick a $j$-dependent sequence
$(x_i,f_i)_{i=1}^{n_{2j-1}}$ with $x_{2i}\in Y$ and $x_{2i-1}\in
Z$ for any $i=1,\dots,n_{2j-1}/2$. Let
$$
y=\frac{m_{2j-1}}{n_{2j-1}}\sum_{i=1}^{n_{2j-1}/2}x_{2i}\in
Y, \ \ \ \
z=\frac{m_{2j-1}}{n_{2j-1}}\sum_{i=1}^{n_{2j-1}/2}x_{2i-1}\in Z
$$
Then by Proposition \ref{dep-est} we have $\|y-z\|\leq 240m_{2j-1}^{-1}$,
whereas
$$
\|y+z\|\geq \frac{1}{m_{2j-1}}\sum_{i=1}^{n_{2j-1}}f_i(y+z)\geq 1
$$
which, as $j$ is arbitrarily big, ends the proof. 
\end{proof}

\begin{proof}[Proof of Theorem \ref{main}]
Apply Proposition \ref{isometric}, Corollary \ref{reflexive} and Theorem \ref{not-complemented}.    
\end{proof}

\begin{remark}
Performing the above construction with the spread $S_{\N,2\N}:c_{00}\ni \sum_{i=1}^\infty a_ie^*_i\mapsto\sum_{i=1}^\infty a_{2i}e^*_i\in c_{00}$ instead of $\Lambda$, we obtain a version of space $X(\mathcal{S})$ of \cite{gm2}, modeled by the proper family of spreads generated by the spread $S_{\N,2\N}$ in the framework of \cite{at}, with $Y=\overline{\textrm{span}}\{e_{2i}:i\in\N\}$ isometric to $X(\mathcal{S})$ and 1-complemented in $X(\mathcal{S})$. 
\end{remark}

\end{document}